Complete the square for this function:
Y=-0.06d^2+2.13

Answer by stanbon(75887)

Complete the square for this function:
Y=-0.06d^2+2.13
-----
-0.06(d-0)^2 = y-2.13
(d-0)^2 = -(100/6)(y-2.13)
-----
(d-0)^2 = -(50/3)(y-2.13)
==============================
Cheers,
Stan H.
=================